OpenStudy (anonymous):

HHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH EEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEELLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPME PLEASE

OpenStudy (anonymous):

Solve for x: |x| − 8 = −5

OpenStudy (akashdeepdeb):

|x| can be = +x or -x So put both the values and get 2 answers. Got it? :)

OpenStudy (anonymous):

yes

OpenStudy (akashdeepdeb):

|x| can be = +x or -x Because !-x| = x = |+x| :D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Sammy224 can you help

OpenStudy (anonymous):

@Hero can you help

OpenStudy (akashdeepdeb):

What? I thought you meant, 'I understood', by the 'yes'. Do you want me to explain? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

|x| - 8 + 8 = - 5 + 8 |x| = 3 or -3

OpenStudy (anonymous):

ty